You bought shares of stock from Tesla, Target, and Toyota. You feel like every year in the future will be randomly either "spectacular" or "depressing".

  

Rate of Return if State Occurs
State of Economy Probability of
State of Economy
Stock of Tesla Stock of Target Stock of Toyota
Spectacular 0.70 0.13 0.05 0.13
Depressing 0.30 0.13 0.13 0.13

  

Requirement 1:

If you bought an equally weighted portfolio of Tesla, Target, and Toyota shares, what return would you expect to receive every year on your stock investment? (Do not round your intermediate calculations.)

  (Click to select)   11.13%   13.63%   23.16%   25.93%   5.40%

  

Requirement 2:

What do you expect the variance of your portfolio returns to be, if you invested 10 percent of your money in Tesla's shares, same proportion in Target's shares, and 80 percent in Toyota's shares? (Do not round your intermediate calculations.)

  (Click to select)   0.008013   0.005513   0.007713   0.003513   0.000013


rev: 09_20_2012, 10_19_2015_QC_CS-21001, 03_30_2016_QC_CS-46530

This is the last question in the assignment. To submit, use Alt + S. To access other questions, proceed to the question map button.Next Visit question map

Question10of10Total10 of 10

Prev
Answer #1

Requirement 1:

E(r) = \sum_{i=1}^{n} [Pi x Ri]

E(rTesla) = [0.70 x 0.13] + [0.30 x 0.13] = 0.091 + 0.039 = 0.13, or 13%

E(rTarget) = [0.70 x 0.05] + [0.30 x 0.13] = 0.035 + 0.039 = 0.074, or 7.40%

E(rToyota) = [0.70 x 0.13] + [0.30 x 0.13] = 0.091 + 0.039 = 0.13, or 13%

E(rP) = [(1/3) x 13%] + [(1/3) x 7.40%] + [(1/3) x 13%] = 4.33% + 2.47% + 4.33% = 11.13%

Hence, Option "A" is correct.

Requirement 2:

E(rP) = [0.10 x 13%] + [0.10 x 7.40%] + [0.80 x 13%] = 1.30% + 0.74% + 10.40%

= 12.44%, or 0.1244

\sigma2 = \sum_{i=1}^{n} [Pi x (E(rP) - Ri)2]

= [0.10 x (0.1244 - 0.13)2] + [0.10 x (0.1244 - 0.074)2] + [0.80 x (0.1244 - 0.13)2]

= 0.000003136 + 0.000254016 + 0.000025088 = 0.00028224